The measure of one angle is 7 times the measure of its complement.Find the measure of each angle. Thanks because i'm really confused on this.....
Found 2 solutions by macston, lwsshak3:
Answer by macston(5194)

The sum of complementary angles is 90 degrees, so:
If the smaller angle is A,
A+7A=90 degrees
8A=90 degrees Divide each side by 8
A=90/8=11.25 degrees ANSWER The smaller angle is 11.25 degrees.
Larger angle=7A=7(11.25)=78.75 degrees ANSWER The larger angle is 78.75 degrees.

Answer by lwsshak3(11628)

The measure of one angle is 7 times the measure of its complement.Find the measure of each angle.
***
let x=angle
its complement=90-x
x=7(90-x)
x=630-7x
8x=630
x=78.75
90-x=11.25
angle=78.75˚
its complement=11.25˚